OpenStudy (anonymous):

How does the intergal of tan2t equal to (1/2)ln abs(sec2t)+c?

OpenStudy (anonymous):

that is the identity, but if you want to know how to get it, you just do tan = sin/cos \[\int tan(2t)dt=\int \dfrac{sin(2t)}{cos(2t)}dt\] Let u = cos (2t) --> du = -2 sin(2t)dt --> sin(2t)dt = -du/2 so, your integral is \[\int -\dfrac{du}{2u} =-1/2 ln|u| +C \] and then -1/2 ln |u| = 1/2ln|u|^(-) , replace u = cos(2t) you have \[1/2ln|(cos(2t)|^{-} =1/2 ln|sec(2t)|\] therefore, your integral = that + C